# arXiv:2603.02011v1[cond-mat.str-el] 02 Mar 2026 thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Introduction
Microscopically, excitonic phenomena can be predicted from the Bethe–Salpeter equation (BSE), which incorporates the single-particle electronic structure and the screened electron–hole interaction [22, 77, 36, 13, 42]. While for periodic systems a momentum space formulation of excitons is successful, for incommensurate modulations, such as those arising in moiré materials, a fully real-space formulation for excitonic effects becomes inevitably necessary. In these cases, conventional methods face severe computational limitations, as even at the single-particle level real-space methods scale poorly with system size [77, 82, 29, 75]. For excitons, the situation is more demanding, since the two-particle Hilbert space scales quadratically with the number of lattice sites. In super-moiré materials, where characteristic length scales can exceed those of standard moiré systems by orders of magnitude, the effective lattice may involve millions or even billions of sites [11, 88], rendering brute-force approaches computationally prohibitive.
@@
Here, we demonstrate a tensor-network (TN) based methodology to access the real-space exciton spectra in super-moire systems. Our methodology relies on combining a real-space TN representation of the electron–hole Hilbert space, and tensor-network kernel polynomial Chebyshev algorithms with a high-resolution spectral kernel. Our approach enables the solution of the real-space BSE on lattices exceeding one billion sites, both in 1D and 2D, surpassing by many orders of magnitude the capabilities of conventional BSE solvers. We show that these techniques allow probing the emergence of exciton moire minibands, together with imaging the microscopic exciton localization, both in quasicrystalline and super-moire systems.

### v2

# thanks: These authors contributed equallythanks: These authors contributed equally

@@ Introduction
Microscopically, excitonic phenomena can be predicted from the Bethe–Salpeter equation (BSE), which incorporates the single-particle electronic structure and the screened electron–hole interaction [23, 79, 37, 14, 44, 41, 8]. While for periodic systems a momentum space formulation of excitons is successful, for incommensurate modulations, such as those arising in moiré materials, a fully real-space formulation for excitonic effects becomes inevitably necessary. In these cases, conventional methods face severe computational limitations, as even at the single-particle level, real-space methods scale poorly with system size [79, 84, 30, 77, 41, 8]. For excitons, the situation is more demanding, since the two-particle Hilbert space scales quadratically with the number of lattice sites. In super-moiré materials, where characteristic length scales can exceed those of standard moiré systems by orders of magnitude, the effective lattice may involve millions or even billions of sites [12, 90], rendering brute-force approaches computationally prohibitive.
@@
Here, we demonstrate a tensor-network (TN) based methodology to access the real-space exciton spectra in super-moire systems. Our methodology relies on combining a real-space TN representation of the electron–hole Hilbert space, and tensor-network kernel polynomial Chebyshev algorithms with a high-resolution spectral kernel. Our approach enables the solution of the real-space BSE on lattices exceeding one billion sites, both in 1D and 2D, surpassing by many orders of magnitude the capabilities of conventional BSE solvers [41, 8]. We show that these techniques allow probing the emergence of exciton moire minibands, together with imaging the microscopic exciton localization, both in quasicrystalline and super-moire systems.